# Conserved Quantities in Expanding Gödel Cosmology

### v3

@@ I Introduction
Recent discoveries by the James Webb Space Telescope (JWST) of an apparent bias in the alignment of the axes of rotation of galaxies throughout the observable universe  JWST has added weight to the earlier evidence of the same phenomenon discovered from, for example, the SDSS data sets  SDSS . One mechanism which has been proposed as driving this alignment is that of a rotating cosmology  JWST . A rotating cosmology has also been suggested as possible method of resolving the Hubble tension between early and late time derived values for the Hubble constant  HUBBLETENSION .
@@
In this work we build on both the previous work in gauge invariant conserved quantities in FRW, for example the curvature perturabtion $\zeta$ , as well as our previous work in Lemaître-Tolman-Bondi (LTB) Bondi and Lemaître spacetimes  Leithes2014 where we construct a similar conserved quantity ${\zeta}_{\rm{SMTP}}$ which was conserved in those spacetimes as well as in FRW.
@@ III.1 Background
where $m$, $\kappa$ and $\sigma$ are simply constants. It can be seen that in the limit $\sigma=m=0$ and $\kappa=1$ we recover the FRW metric. For a more detailed description of the origin and form of these constants see  REFEG . The off-diagonal components of the background metric will clearly lead to more complicated equations in the following sections than those employed in FRW and to some degree even LTB and Lemaître spacetimes.
@@
In order to construct the Einstein equations we must first find the connections, zero and non-zero, denoted by the Christoffel symbols below. In  REFEG it seems that gauge selections have been made to yield a remarkably small number of non-zero connections - seven including duplicates. Here we make no initial gauge choices choosing to leave our expressions fully general, only later fixing gauges when dealing with perturbations in order to find gauge invariant conserved quantities. As such we find,
[…]
with the remaining thirty-eight connections all being zero. Here the “prime” denotes the derivative with respect to the spatial coordinate $x$.
@@
From Eq. (1) we find the Einstein equations are, for the 000-0 component,
For the 010-1 component we find,
For the 020-2 component we find,
For the 111-1 component we find,
For the 121-2 component we find,
For the 222-2 component we find,
For the 333-3 component we find,
The 030-3, 131-3 and 232-3 components are all identically zero.
@@ Appendix A Additional Material for Background EG Spacetime
A.1 The Ricci Tensor Components

### v4

@@ I Introduction
Recent discoveries by the James Webb Space Telescope (JWST) of an apparent bias in the alignment of the axes of rotation of galaxies throughout the observable universe  JWST (specifically that 50% more galaxies surveyed rotate in the opposite direction relative to the Milky Way galaxy compared to those rotating in the same direction) has added weight to the earlier evidence of the same phenomenon discovered from, for example, the SDSS data sets  SDSS . One mechanism which has been proposed as driving this alignment is that of a rotating cosmology  JWST . Specifically in  JWST it is suggested that a cosmological-scale axis of alignment for galaxy rotations implies that the entire cosmology may intrinsically have such an axis, as would be the case in a rotating cosmology.
A rotating cosmology has also been suggested as possible method of resolving the Hubble tension between early and late time derived values for the Hubble constant  HUBBLETENSION .
@@
In this work we build on both the previous work in gauge invariant conserved quantities in FRW, for example the curvature perturabtion $\zeta$ , as well as our previous work in Lemaître-Tolman-Bondi (LTB) Bondi and Lemaître spacetimes  Leithes2014 where we construct a similar conserved quantity ${\zeta}_{\rm{SMTP}}$ which was conserved in those spacetimes as well as in FRW. To avoid confusion we point out as this stage that FRW is the standard model used in cosmology today and has the background metric,
$ds^{2}=-dt^{2}+a(t)^{2}\delta_{ij}dx^{i}dx^{j}\,$, (1)
where $a(t)$ is the time only dependent scale factor. LTB spacetime has the following background metric,
$ds^{2}=-dt^{2}+X^{2}(r,t)dr^{2}+Y^{2}(r,t)\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (2)
where $X(r,t)$ and $Y(r,t)$ are scale factors dependent upon both the radial spatial and time co-ordinates and are not independent. Finally, Lemaître spacetime differs from LTB spacetime and has the following background metric,
$ds^{2}=-f^{2}dt^{2}+X^{2}(r,t)dr^{2}+Y^{2}(r,t)\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)\,$, (3)
where $f$ is an additional third scale factor, $f\equiv f(r,t)$. These metrics are only shown here for reference and for comparison with the EG metric to follow. For a more detailed analysis of these other cosmologies see e.g. Leithes2014 .
@@ III.1 Background
where $m$, $\kappa$ and $\sigma$ are simply constants. It can be seen that in the limit $\sigma=m=0$ and $\kappa=1$ we recover the FRW metric. For a more detailed description of the origin and form of these constants see  REFEG .
It should be noted here that from the form of $\omega$ in Eq. (16) and the term $\frac{m}{2a}$ it can easily be seen that with the rapid expansion of the universe and commensurate increase in the scale factor, $a$, the magnitude of the rotation, $\omega$, rapidly decreases to negligible levels in late times as expected. Naturally, angular momentum is conserved.
The off-diagonal components of the background metric will clearly lead to more complicated equations in the following sections than those employed in FRW and to some degree even LTB and Lemaître spacetimes.
@@
In order to construct the Einstein equations we must first find the connections, zero and non-zero, denoted by the Christoffel symbols below. In  REFEG it seems that gauge selections have been made to yield a remarkably small number of non-zero connections - seven including duplicates. Here we make no initial gauge choices choosing to leave our expressions fully general, only later fixing gauges when dealing with perturbations in order to find gauge invariant conserved quantities. As such we find twenty-six non-zero connections as shown in Appendix A.1, with the remaining thirty-eight connections all being zero. Here a “prime” denotes the derivative with respect to the spatial coordinate $x$.
@@
From Eq. (4) we find the Einstein equations are as given in Appendix A.4.
@@ Appendix A Additional Material for Background EG Spacetime
A.1 The Non-zero Christoffel Symbols
The non-zero connections are,
[…]
A.2 The Ricci Tensor Components
@@ A.3 The Ricci Scalar
A.4 The Einstein Field Equations
For the 000-0 component we find,
For the 010-1 component we find,
For the 020-2 component we find,
For the 111-1 component we find,
For the 121-2 component we find,
For the 222-2 component we find,
For the 333-3 component we find,
The 030-3, 131-3 and 232-3 components are all identically zero.